# Regret Analysis for Randomized Gaussian Process Upper Confidence Bound

### v1

@@ 5.1 Synthetic Function Experiment
We use synthetic functions generated from ${\cal G}{\cal P}(0,k)$, where $k$ is Gaussian kernel with a length scale parameter $\ell=0.1$ and the input dimension $d=3$. We set the noise variance $\sigma^{2}=10^{-4}$. The input domain consists of equally divided points in $[0,0.9]$, that is, ${\cal X}=\{0,0.1,\dots,0.9\}^{d}$ and $|{\cal X}|=1000$. Figure 1(a) shows the theoretical confidence parameters of GP-UCB, RGP-UCB with ${\rm Gamma}(\kappa_{t}=\log(|{\cal X}|t^{2})/\log(1.5),\theta=1)$, and IRGP-UCB with a two-parameter exponential distribution. We can see that the confidence parameters for GP-UCB and RGP-UCB are considerably large due to logarithmic increase, particularly in the later iterations. In contrast, the confidence parameter of IRGP-UCB is drastically smaller than that of GP-UCB and not changed with respect to the iteration. Therefore, we can observe that IRGP-UCB alleviates the over-exploration.
Next, we report the simple regret in Figure 1(b). Ten synthetic functions and ten initial training datasets are randomly generated. Thus, the average and standard error for $10\times 10$ trials are reported. The hyperparameters for GP are fixed to those used to generate the synthetic functions. The confidence parameters for GP-UCB, RGP-UCB, and IRGP-UCB are set as in Figure 1(a). We can see that the regrets of GP-UCB, RGP-UCB, TS, and JES converge slowly. We conjecture that this slow convergence came from over-exploration. On the other hand, EI, PIMS, MES, and IRGP-UCB show fast convergence. In particular, IRGP-UCB achieves the best average in most iterations. These results suggest that IRGP-UCB bridges the gap between theory and practice in contrast to GP-UCB and RGP-UCB.
(a) Confidence parameter.
(b) Simple regrets.
Figure 1: Figure (a) shows the confidence parameters of GP-UCB-based methods. For GP-UCB, the solid line represents $\beta_{t}$. For RGP-UCB and IRGP-UCB, the solid lines and shaded area represent the expectations $\mathbb{E}[\zeta_{t}]$ and $95\%$ credible intervals, respectively. Figure (b) shows the average and standard errors of simple regrets.
@@ 5.2 Benchmark Function Experiments
(a) Benchmark functions.
(b) Materials datasets.
Figure 2: Average and standard errors of simple regret.
We employ three benchmark functions called Holder table ($d=2$), Cross in tray($d=2$), and Ackley ($d=4$) functions, whose analytical forms are shown at https://www.sfu.ca/~ssurjano/optimization.html. For each function, we report the average and standard error for $10$ trails using ten random initial datasets ${\cal D}_{0}$, where $|{\cal D}_{0}|=2^{d}$. We set the noise variance $\sigma^{2}=10^{-4}$. We used the Gaussian kernel with automatic relevance determination, whose hyperparameter was selected by marginal likelihood maximization per $5$ iterations (Rasmussen and Williams, 2005).
In this experiment, since the input domain is continuous, the theoretical choice of the confidence parameter contains an unknown variable. Thus, we use the heuristic choice for confidence parameters. For GP-UCB, we set the confidence parameter as $\beta_{t}=0.2d\log(2t)$, which is the heuristics used in (Kandasamy et al., 2015, 2017). For RGP-UCB, we set $\zeta_{t}\sim{\rm Gamma}(\kappa_{t},\theta=1)$ with $\kappa_{t}=0.2d\log(2t)$ since $\mathbb{E}[\zeta_{t}]$ must have the same order as $\beta_{t}$ (note that $\mathbb{E}[\zeta_{t}]=\theta\kappa_{t}$). For IRGP-UCB, we set $s=d/2$ and $\lambda=1/2$. Note that $\lambda$ is equal to the theoretical setting, and $s$ captures the dependence on $d$, as shown in Corollary 4.2.
Figure 2(a) shows the results. In the Holder table function, JES shows faster convergence until 40 iterations. However, the regrets of all baseline methods stagnate in $[10^{-1},10^{0}]$. In contrast, the regret of IRGP-UCB converges to $10^{-3}$ until 60 iterations. In the Cross in tray function, IRGP-UCB showed rapid convergence at approximately 50 iterations. In the Ackley function, IRGP-UCB constantly belonged to the top group and showed minimum regret after 125 iterations. Since the input dimension is relatively large, TS and JES, which depend on the sampled maxima $\boldsymbol{x}_{*}$, deteriorate by over-exploration. Throughout the experiments, IRGP-UCB outperforms TS, PIMS, GP-UCB, and RGP-UCB, which achieves sub-linear BCR bounds, and EI and MES, which are practically well-used. This result supports the effectiveness of randomization using the two-parameter exponential distribution.
@@ 5.3 Real-World Dataset Experiments
This section provides the experimental results on the materials datasets provided in (Liang et al., 2021). In the perovskite dataset (Sun et al., 2021), we optimize environmental stability with respect to composition parameters for halide perovskite ($d=3$ and $|{\cal X}|=94$). In the P3HT/CNT dataset (Bash et al., 2021), we optimize electrical conductivity with respect to composition parameters for carbon nanotube polymer blend ($d=5$ and $|{\cal X}|=178$). In the AgNP dataset (Mekki-Berrada et al., 2021), we optimize the absorbance spectrum of synthesized silver nanoparticles with respect to processing parameters for synthesizing triangular nanoprisms ($d=5$ and $|{\cal X}|=164$). See (Liang et al., 2021) for more details about each dataset.
@@
Figure 2(b) shows the results. In the perovskite dataset, IRGP-UCB constantly belonged to the top group and showed the best performance after 20 iterations. In the P3HT/CNT dataset, EI converged to the smallest value after 60 iterations. On the other hand, IRGP-UCB shows faster convergence during the first 20 iterations. In the AgNP dataset, IRGP-UCB found the optimal point until 42 iterations in all the trials earliest. In this experiment, heuristic methods, EI, MES, and JES, showed worse performance and required at least 60 function evaluations to find the optimal point. Consequently, we can observe the effectiveness of IRGP-UCB against real-world datasets.
@@ 6 Conclusion
This paper analyses the expected regret of the IRGP-UCB algorithm in the Bayesian setting. In addition, to fairly compare with the deterministic algorithm, we analyze the conditional expected regret. In both analyses, if the input domain is finite, increasing the confidence parameter is avoided, and the regret upper bound is slightly tightened. Finally, we show the practical effectiveness of the IRGP-UCB through the numerical experiments, including the application to the materials dataset.
Several directions for future works can be considered. First, whether we can show the high probability bounds of regrets without increasing the confidence parameter is of interest. Second, the tighter BCR bounds for the continuous domain are also an important direction. Last, as the randomized LST has been proposed (Inatsu et al., 2024b), we may be able to extend IRGP-UCB to the other various practical settings, where the usual UCB-based methods have been extended, for example, multi-objective BO (Paria et al., 2020; Zuluaga et al., 2016; Suzuki et al., 2020; Inatsu et al., 2024a), multi-fidelity BO (Kandasamy et al., 2016, 2017; Takeno et al., 2020, 2022a), parallel BO (Contal et al., 2013; Desautels et al., 2014), high-dimensional BO (Kandasamy et al., 2015), cascade BO (Kusakawa et al., 2022), and robust BO(Bogunovic et al., 2018; Iwazaki et al., 2021; Inatsu et al., 2022).
@@ Acknowledgements
This work was supported by MEXT KAKENHI 21H03498, 22H00300, MEXT Program: Data Creation and Utilization-Type Material Research and Development Project Grant Number JPMXP1122712807, JSPS KAKENHI Grant Number JP20H00601, JP21J14673, JP23K16943, JP23K19967 and JP24K20847, and JST ACT-X Grant Number JPMJAX23CD, RIKEN Center for Advanced Intelligence Project.
@@ Appendix A Proofs for Expected Regret Bounds of IRGP-UCB
The main difference between Lemma 4.1 and (Lemma 5.1 in Srinivas et al., 2010) is that Lemma 4.1 does not consider the union bound for all $t\geq 1$. That is, Lemma 5.1 in Srinivas et al. (2010) considers the following probability:
@@
where we use an elementary result of Gaussian distribution shown in Lemma D.3 in Appendix D. Then, we can obtain the following bound:
@@
Using Lemma 4.2, we can obtain the proof of the following theorem for the finite input domain:
@@
Purely for the sake of analysis, we used a set of discretization ${\cal X}_{t}\subset{\cal X}$ for $t\geq 1$. For any $t\geq 1$, let ${\cal X}_{t}\subset{\cal X}$ be a finite set with each dimension equally divided into $\tau_{t}=bdrt^{2}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$. Thus, $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we define $[\boldsymbol{x}]_{t}$ as the nearest points in ${\cal X}_{t}$ of $\boldsymbol{x}\in{\cal X}$.
@@
First, we consider the first term $\sum_{t=1}^{T}\mathbb{E}\bigl{[}f(\boldsymbol{x}^{*})-f([\boldsymbol{x}^{*}]_{t})\bigr{]}$. From Lemma D.2, we can obtain the following:
@@
where $u_{t}$ in Lemma D.2 corresponds to $t^{2}$.
@@ Appendix B Proofs for Conditional Expected Regret Bounds of IRGP-UCB
The definition of the subgaussian random variable is shown below (Definition 3.5 of van Handel, 2016):
@@
Several equivalent definitions are known (Proposition 2.5.2 of Vershynin, 2018). In particular, we use the following relations:
@@
for all $c\geq 0$ and $\mathbb{E}[X]=0$. Then, $X$ is at least $9A^{2}$-subgaussian.
@@
For example, by combining Problems 3.1 (d) and (e) of van Handel (2016), we can obtain this result. ∎
@@
See, for example, Problem 3.1 (c) of van Handel (2016) or the proof for (v) $\Rightarrow$ (i) of Proposition 2.5.2 in (Vershynin, 2018). ∎
Furthermore, we use the following Azuma’s inequality (Lemma 3.7 of van Handel, 2016; Azuma, 1967):
@@
To show the subgaussian property, we use the following proposition shown in, for example, Theorem 5.2.2 of Vershynin (2018), Proposition 4 of Zeitouni (2014), and Lemma 6.4 of Biskup (2020):
@@
$\leq 6\sqrt{T\log(2/\delta)}+\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log(2/\delta)}+2\log(2/\delta)\right)}$
$=O\left(\sqrt{T\gamma_{T}\log|{\cal X}|+\gamma_{T}\sqrt{T\log(1/\delta)}}+\gamma_{T}\log(1/\delta)\right),$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
For any $\{\zeta_{t}\}_{t\geq 1}$, we obtain
@@
First, regarding $A_{2}$, the following inequality holds with probability at least $21-\delta/$ 2,
@@
where the last inequality holds with high probability. Note that, regarding the first inequality, $\sqrt{\sum_{t=1}^{T}\zeta_{t}}$ can be put outside of the conditional expectation since the expectation is taken with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$.
@@
In addition, from the definition of mearuarability, $\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid{\cal H}_{t}\right]$ is ${\cal H}_{t}$-mesuarable.
@@
Note that for any fixed ${\cal H}_{t-1}$, $h(\boldsymbol{X})\mid{\cal H}_{t-1}\overset{\mathrm{d}}{=}\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid\zeta_{t}\right]\mid{\cal H}_{t-1}$, where $\boldsymbol{X}\sim{\cal N}(\boldsymbol{0},\boldsymbol{I}_{2})$ and =$\overset{\mathrm{d}}{=}$ means the equivalence of the distribution, since $\|\boldsymbol{X}\|_{2}^{2}$ follows the exponential distribution with $\lambda=1/2$. Then, for any fixed ${\cal H}_{t-1}$, $h$ is $1$-Lipschitz function as shown below:
@@
The inequalities (a), (b), and (c) are derived by (a) $\sigma_{t-1}(\boldsymbol{x})\leq 1$ and that the random variables depending on ${\cal D}_{t-1}$ is removed, (b) the monotone decreasing property of the gradient of $\sqrt{\cdot}$, and (c) the triangle inequality. Therefore, from Proposition B.3 and $h(\boldsymbol{X})\mid{\cal H}_{t-1}\overset{\mathrm{d}}{=}\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid\zeta_{t}\right]\mid{\cal H}_{t-1}$,
@@
Finally, since we confirm that $A_{1}$ is the sum of random variables that satisfy the condition of Azuma’s inequality, $A_{1}$ is $18T$-subgaussian. Therefore, from Proposition B.2, the following inequality holds with probability at least $21-\delta/$ 2,
$A_{1}\leq 6\sqrt{T\log(2/\delta)}.$
Hence, using the union bound for the bounds of $A_{1}$ and $A_{2}$, we obtain the desired result. ∎
@@
$\leq\frac{\pi^{2}}{6}+6\sqrt{T\log(2/\delta)}+\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log(2/\delta)}+2\log(2/\delta)\right)},$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
Purely for the sake of analysis, we used a set of discretization ${\cal X}_{t}\subset{\cal X}$ for $t\geq 1$. For any $t\geq 1$, let ${\cal X}_{t}\subset{\cal X}$ be a finite set with each dimension equally divided into $\tau_{t}=bdrt^{2}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$. Thus, $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we define $[\boldsymbol{x}]_{t}$ as the nearest points in ${\cal X}_{t}$ of $\boldsymbol{x}\in{\cal X}$.
@@
Then, the first term can be bounded from above by $\pi^{2}/6$ as same as the proof of Theorem 4.2. For the second term, we see that
@@
Then, remained proof is almost the same as Theorem 4.3. ∎
Appendix C Proofs for High Probability Regret Bounds of IRGP-UCB
C.1 Proof of Theorem 4.5

### v3

@@ 5 Experiments
Figure 1: This figure shows the confidence parameters of GP-UCB-based methods. For GP-UCB, the solid line represents $\beta_{t}$. For RGP-UCB and IRGP-UCB, the solid lines and shaded areas represent the expectations $\mathbb{E}[\zeta_{t}]$ and $95\%$ credible intervals, respectively.
Figure 2: This figure shows the average and standard errors of simple regrets.
@@ 5.1 Synthetic Function Experiment
We use synthetic functions generated from ${\cal G}{\cal P}(0,k)$, where $k$ is a Gaussian kernel with a length scale parameter $\ell=0.1$ and the input dimension $d=3$. We set the noise variance $\sigma^{2}=10^{-4}$. The input domain consists of equally divided points in $[0,0.9]$, that is, ${\cal X}=\{0,0.1,\dots,0.9\}^{d}$ and $|{\cal X}|=1000$. Figure 1 shows the theoretical confidence parameters of GP-UCB, RGP-UCB with ${\rm Gamma}(\kappa_{t}=\log(|{\cal X}|t^{2})/\log(1.5),\theta=1)$, and IRGP-UCB with a two-parameter exponential distribution. We can see that the confidence parameters for GP-UCB and RGP-UCB are considerably large due to a logarithmic increase, particularly in the later iterations. In contrast, the confidence parameter of IRGP-UCB is drastically smaller than that of GP-UCB and does not change with respect to the iteration. Therefore, we can observe that IRGP-UCB alleviates the over-exploration.
Next, we report the simple regret in Figure 2. Ten synthetic functions and ten initial training datasets are randomly generated. Thus, the average and standard error for $10\times 10$ trials are reported. The hyperparameters for GP are fixed to those used to generate the synthetic functions. The confidence parameters for GP-UCB, RGP-UCB, and IRGP-UCB are set as in Figure 1. We can see that the regrets of GP-UCB, RGP-UCB, TS, and JES converge slowly. We conjecture that this slow convergence came from over-exploration. On the other hand, EI, PIMS, MES, and IRGP-UCB show fast convergence. In particular, IRGP-UCB achieves the best average in most iterations. These results suggest that IRGP-UCB bridges the gap between theory and practice in contrast to GP-UCB and RGP-UCB.
Figure 3: Average and standard errors of simple regret on benchmark functions.
@@ 5.2 Benchmark Function Experiments
We employ three benchmark functions called Holder table ($d=2$), Cross in tray($d=2$), and Ackley ($d=4$) functions, whose analytical forms are shown at https://www.sfu.ca/~ssurjano/optimization.html. For each function, we report the average and standard error for $10$ trials using ten random initial datasets ${\cal D}_{0}$, where $|{\cal D}_{0}|=2^{d}$. We set the noise variance $\sigma^{2}=10^{-4}$. We used the Gaussian kernel with automatic relevance determination, whose hyperparameter was selected by marginal likelihood maximization per $5$ iterations (Rasmussen and Williams, 2005).
In this experiment, since the input domain is continuous, the theoretical choice of the confidence parameter contains an unknown variable. Thus, we use the heuristic choice for confidence parameters. For GP-UCB, we set the confidence parameter as $\beta_{t}=0.2d\log(2t)$, which is the heuristic used in (Kandasamy et al., 2015, 2017). For RGP-UCB, we set $\zeta_{t}\sim{\rm Gamma}(\kappa_{t},\theta=1)$ with $\kappa_{t}=0.2d\log(2t)$ since $\mathbb{E}[\zeta_{t}]$ must have the same order as $\beta_{t}$ (note that $\mathbb{E}[\zeta_{t}]=\theta\kappa_{t}$). For IRGP-UCB, we set $s=d/2$ and $\lambda=1/2$. Note that $\lambda$ equals the theoretical setting, and $s$ captures the dependence on $d$, as shown in Corollary 4.2.
Figure 3 shows the results. In the Holder table function, JES shows faster convergence until 40 iterations. However, the regrets of all baseline methods stagnate in $[10^{-1},10^{0}]$. In contrast, the regret of IRGP-UCB converges to $10^{-3}$ until 60 iterations. In the Cross in tray function, IRGP-UCB showed rapid convergence at approximately 50 iterations. In the Ackley function, IRGP-UCB constantly belonged to the top group and showed minimum regret after 125 iterations. Since the input dimension is relatively large, TS and JES, which depend on the sampled maxima $\boldsymbol{x}_{*}$, deteriorate by over-exploration. Throughout the experiments, IRGP-UCB outperforms TS, PIMS, GP-UCB, and RGP-UCB, which achieve sub-linear BCR bounds, and EI and MES, which are practically well-used. This result supports the effectiveness of randomization using the two-parameter exponential distribution.
Figure 4: Average and standard errors of simple regret on real-world datasets.
@@ 5.3 Real-World Dataset Experiments
This section provides the experimental results on the materials datasets provided in (Liang et al., 2021). In the perovskite dataset (Sun et al., 2021), we optimize environmental stability with respect to composition parameters for halide perovskite ($d=3$ and $|{\cal X}|=94$). In the P3HT/CNT dataset (Bash et al., 2021), we optimize electrical conductivity with respect to composition parameters for the carbon nanotube polymer blend ($d=5$ and $|{\cal X}|=178$). In the AgNP dataset (Mekki-Berrada et al., 2021), we optimize the absorbance spectrum of synthesized silver nanoparticles with respect to processing parameters for synthesizing triangular nanoprisms ($d=5$ and $|{\cal X}|=164$). See (Liang et al., 2021) for more details about each dataset.
@@
Figure 4 shows the results. In the perovskite dataset, IRGP-UCB constantly belonged to the top group and showed the best performance after 20 iterations. In the P3HT/CNT dataset, EI converged to the smallest value after 60 iterations. On the other hand, IRGP-UCB shows faster convergence during the first 20 iterations. In the AgNP dataset, IRGP-UCB found the optimal point after 42 iterations in all the trials. In this experiment, heuristic methods, EI, MES, and JES, showed worse performance and required at least 60 function evaluations to find the optimal point. Consequently, we can observe the effectiveness of IRGP-UCB against real-world datasets.
@@ 6 Conclusion
This paper analyzes the expected regret of the randomized variant of GP-UCB called IRGP-UCB in the Bayesian setting. Moreover, to fairly compare with the deterministic algorithm, we obtain the high-probability upper bound of the conditional expected regret averaged over the randomness of the GP and the noise, and conditioned on the algorithm’s randomness. In both analyses, if the input domain is finite, increasing the confidence parameter, which can harm the practical efficiency, is avoided. In addition, by showing the specific instance where GP-UCB using a constant confidence parameter incurs linear expected cumulative regret, we confirmed that randomization plays a key role in avoiding increasing confidence parameters. Finally, we show the practical effectiveness of the IRGP-UCB through the numerical experiments, including the application to the materials dataset.
Several directions for future work can be considered. First, whether we can show high-probability regret bounds without increasing the confidence parameter is of interest. Second, avoiding an increase in the confidence parameter for the continuous domain is also an important direction. Last, as with the randomized LSE proposed recently (Inatsu et al., 2024b), we may be able to extend IRGP-UCB to various other practical settings, where the usual GP-UCB-based methods have been extended, for example, multi-objective BO (Paria et al., 2020; Zuluaga et al., 2016; Suzuki et al., 2020; Inatsu et al., 2024a), multi-fidelity BO (Kandasamy et al., 2016, 2017; Takeno et al., 2020, 2022a), parallel BO (Contal et al., 2013; Desautels et al., 2014), high-dimensional BO (Kandasamy et al., 2015), cascade BO (Kusakawa et al., 2022), and robust BO (Bogunovic et al., 2018; Iwazaki et al., 2021; Inatsu et al., 2022).
@@ Acknowledgements
This work was supported by JSPS/MEXT KAKENHI Grant Number (JP20H00601, JP21H03498, 22H00300, JP21J14673, JP23K16943, JP23K19967, and JP24K20847), MEXT Program: Data Creation and Utilization-Type Material Research and Development Project Grant Number JPMXP1122712807, JST ACT-X Grant Number (JPMJAX23CD and JPMJAX24C3), JST PRESTO Grant Number JPMJPR24J6, and RIKEN Center for Advanced Intelligence Project.
@@ Appendix A Proofs for Expected Regret Bounds of IRGP-UCB
The main difference between Lemma 4.1 and [Lemma 5.1 in Srinivas et al., 2010] is that Lemma 4.1 does not consider the union bound for all $t\geq 1$. That is, Lemma 5.1 in Srinivas et al. [2010] considers the following probability:
@@
where we use an elementary result of the Gaussian distribution shown in Lemma E.3 in Appendix E. Then, we can obtain the following bound:
@@
Using Lemma 4.2, we can obtain the proof of the following theorem for finite input domains:
@@
We consider a finite discretized input set ${\cal X}_{t}\subset{\cal X}$ for $t\geq 1$ for theoretical analysis, though ${\cal X}_{t}$ is not related to the actual algorithm. Let ${\cal X}_{t}=\{\frac{r}{\tau_{t}+1},\dots,\frac{\tau_{t}r}{\tau_{t}+1}\}^{d}\subset{\cal X}$, where each coordinate takes values from a uniform grid of $\tau_{t}=bdru_{t}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$ points. Thus, $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we define $[\boldsymbol{x}]_{t}$ as the nearest point in ${\cal X}_{t}$ of $\boldsymbol{x}\in{\cal X}$.
@@
First, we consider the first term $\sum_{t=1}^{T}\mathbb{E}\bigl{[}f(\boldsymbol{x}^{*})-f([\boldsymbol{x}^{*}]_{t})\bigr{]}$. From Lemma E.2, we can obtain the following:
@@
where $u_{t}$ in Lemma E.2 corresponds to $t^{2}$.
@@ Appendix B Proofs for Conditional Expected Regret Bounds of IRGP-UCB
The definition of the subgaussian random variable is shown below [Definition 3.5 of Handel, 2016]:
@@
Several equivalent definitions are known [Proposition 2.5.2 of Vershynin, 2018]. In particular, we use the following relations:
@@
for all $c\geq 0$ and $\mathbb{E}[X]=0$. Then, $X$ is $9A^{2}$-subgaussian.
@@
For example, by combining Problems 3.1 (d) and (e) of Handel [2016], we can obtain this result. ∎
@@
See, for example, Problem 3.1 (c) of Handel [2016] or the proof for (v) $\Rightarrow$ (i) of Proposition 2.5.2 in [Vershynin, 2018]. ∎
Furthermore, we use the following form of Azuma’s inequality [Lemma 3.7 of Handel, 2016, Azuma, 1967]:
@@
To show the subgaussian property the variable $A_{1}$ in the following proof of Theorem 4.3, we use the following proposition shown in, for example, Theorem 5.2.2 of Vershynin [2018], Proposition 4 of Zeitouni [2014], and Lemma 6.4 of Biskup [2020]:
@@
$\Pr\left(\forall T\geq 1,\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}\right]\leq U\bigl{(}T,\delta\bigr{)}\right)\geq 1-\delta,$
where
$U(T,\delta)$ $=6\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)}$
and $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
First, we show an upper bound for some fixed $T\geq 1$. For any $\{\zeta_{t}\}_{t\geq 1}$, we obtain
@@
$=\underbrace{\sum_{t=1}^{T}\biggl{\{}\mathbb{E}_{{\cal D}_{t-1},\zeta_{t}}\left[v_{t}(\boldsymbol{x}_{t})\mid\{\zeta_{i}\}_{i<t}\right]-\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid\{\zeta_{i}\}_{i\leq t}\right]\biggr{\}}}_{A_{1}}+\underbrace{\mathbb{E}\left[\sum_{t=1}^{T}\zeta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t>1}\right]}_{A_{2}},$
@@
First, regarding $A_{2}$, the following inequality holds with probability at least $1-\delta_{2}$,
@@
where the last inequality holds with high-probability since $\sum_{t=1}^{T}\{\zeta_{t}-s_{t}\}$ follows a chi-square distribution with $T$ degrees of freedom. Note that, regarding the first inequality, $\sqrt{\sum_{t=1}^{T}\zeta_{t}}$ can be put outside of the conditional expectation since the expectation is taken with $f$ and $\{\epsilon_{t}\}_{t\geq 1}$.
@@
In addition, from the definition of measurability, $\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid{\cal H}_{t}\right]$ is ${\cal H}_{t}$-mesuarable.
@@
Note that $\bigl{(}h(\boldsymbol{X})\mid{\cal H}_{t-1}\bigr{)}\overset{\mathrm{d}}{=}\bigl{(}\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid\zeta_{t}\right]\mid{\cal H}_{t-1}\bigr{)}$ for any fixed ${\cal H}_{t-1}$, where $\boldsymbol{X}\sim{\cal N}(\boldsymbol{0},\boldsymbol{I}_{2})$ and =$\overset{\mathrm{d}}{=}$ means the equivalence of the distribution, since $\|\boldsymbol{X}\|_{2}^{2}$ follows the exponential distribution with $\lambda=1/2$. Then, for any fixed ${\cal H}_{t-1}$, $h$ is $1$-Lipschitz function as shown below:
@@
The inequalities (a), (b), and (c) are derived by (a) $\sigma_{t-1}(\boldsymbol{x})\leq 1$, (b) the monotone decreasing property of the gradient of $\sqrt{\cdot}$, and (c) the triangle inequality. Therefore, from Proposition B.3 and $\bigl{(}h(\boldsymbol{X})\mid{\cal H}_{t-1}\bigr{)}\overset{\mathrm{d}}{=}\bigl{(}\mathbb{E}_{{\cal D}_{t-1}}\left[v_{t}(\boldsymbol{x}_{t})\mid\zeta_{t}\right]\mid{\cal H}_{t-1}\bigr{)}$,
@@
Finally, since we see that $A_{1}$ is the sum of random variables that satisfy the condition of Azuma’s inequality, $A_{1}$ is $18T$-subgaussian. Therefore, from Proposition B.2, the following inequality holds with probability at least $1-\delta_{1}$,
$A_{1}\leq 6\sqrt{T\log(1/\delta_{1})}.$
Hence, for some fixed $T\geq 1$, we obtain the high-probability bounds of $A_{1}$ and $A_{2}$. Then, substituting $6\delta/(\pi^{2}T^{2})$ into $\delta_{1}$ and $\delta_{2}$ and taking the union bounds for all $T\geq 1$, we conclude the proof. ∎
@@
$\Pr\left(\forall T\geq 1,\mathbb{E}_{f,\{\epsilon_{t}\}_{t\geq 1}}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid\{\zeta_{t}\}_{t\geq 1}\right]\leq U\bigl{(}T,\delta\bigr{)}\right)\geq 1-\delta,$
where
$U(T,\delta)$ $=\frac{\pi^{2}}{6}+6\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)},$
and $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
We consider a finite discretized input set ${\cal X}_{t}\subset{\cal X}$ for $t\geq 1$ for theoretical analysis, though ${\cal X}_{t}$ is not related to the actual algorithm. Let ${\cal X}_{t}=\{\frac{r}{\tau_{t}+1},\dots,\frac{\tau_{t}r}{\tau_{t}+1}\}^{d}\subset{\cal X}$, where each coordinate takes values from a uniform grid of $\tau_{t}=bdru_{t}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$ points. Thus, $|{\cal X}_{t}|=\tau_{t}^{d}$. In addition, we define $[\boldsymbol{x}]_{t}$ as the nearest point in ${\cal X}_{t}$ of $\boldsymbol{x}\in{\cal X}$.
@@
Then, the first term can be bounded from above by $\pi^{2}/6$ as in the proof of Theorem 4.2. For the second term, we see that
@@
Then, the rest of the proof is almost the same as Theorem 4.3. ∎
Appendix C Proofs for High-Probability Regret Bounds of IRGP-UCB